OpenStudy (anonymous):

What is the value of x in the inequality 4x + 6 < 14? x < 2 x < 4 x > 2 x > 4

OpenStudy (saifoo.khan):

x < 2

OpenStudy (anonymous):

Which expression is true? Select answer A (9 - 4) - 3 = 9 - (4 - 3) Select answer B (36 ÷ 4) ÷ 3 = 36 ÷ (4 ÷ 3) Select answer C 9 - (4 - 3) = (9 - 4) - 3 Select answer D (9 × 4) × 3 = 9 × (4 × 3)